OpenStudy (anonymous):

Can someone please help me figure this out or know a formula to use? Jacob purchased 110 shares of stock in a computer company at $92.63 per share and Rae purchased 45 shares of stock in a different computer company for $87.45 per share. After holding the stock for two years, Jacob sold his for a total of $11,120.53, and Rae sold hers for a total of $4,192.10. Which person had a higher ROI and by how much? Jacob: 9.1% more Rae: 6.5% more Jacob: 2.6% more Rae: 2.6% more

ganeshie8 (ganeshie8):

ROI means Return On Investment

ganeshie8 (ganeshie8):

ROI = \(\large \frac{\textbf{final value - initial value}}{\textbf{initial value}}\)

ganeshie8 (ganeshie8):

find the ROI of each separately and see the difference

OpenStudy (anonymous):

Thank you so much!

ganeshie8 (ganeshie8):

np :) wat do u get as answer ?

OpenStudy (anonymous):

I haven't done it yet, I am right now

OpenStudy (anonymous):

I don't get the initial value?

ganeshie8 (ganeshie8):

initial value = (number of shares) * (price of share)

OpenStudy (anonymous):

I got 9.1% more for Jacob

OpenStudy (anonymous):

Could you help me with one more? @ganeshie8

ganeshie8 (ganeshie8):

hey doesnt look correct @surferbb9

ganeshie8 (ganeshie8):

Jacob : initial = 110*92.63 final = 11120.53 ROI = (11120.53 - 110*92.63)/(110*92.63)*100 = 9.1 % Rae : initial : 45*87.45 finanl = 4192.1 ROI = (4192.1-45*87.45)/(45*87.45) = 6.5 %

ganeshie8 (ganeshie8):

difference = 9.1 - 6.5

ganeshie8 (ganeshie8):

= 2.6 % more for Jacob